OpenStudy (anonymous):

Consider the set of families with exactly 4 children. If P (chi;d is a boy)=1/2, find the probablility that one of these families, picked at random, has 4 boys.

Parth (parthkohli):

We basically multiply the probabilities if we want four things to happen at once. \( \color{Black}{\Rightarrow \Large {1 \over 2} \times {1 \over 2} \times{1 \over 2}\times{1 \over 2} }\)

OpenStudy (anonymous):

thanks